# Analog Computing for Signal Processing and Communications – Part I: Computing with Microwave Networks

### v2

@@ V-B Computing the Kalman Filter with a MiLAC
The Kalman filter in Alg. 4 can be computed by performing the most computationally expensive steps in the analog domain through a MiLAC, as shown in Alg. 5. Note that Alg. 5 takes in input the matrix inverse $\mathbf{R}_{t-1\mid t-1}^{-1}$ and return the matrix inverse $\mathbf{R}_{t\mid t}^{-1}$, differently from Alg. 4, but without altering the logic of the Kalman filter. In the “prediction” phase, Step 1 in Alg. 4 has low computational complexity, i.e., $8X^{2}$ real operations (see Appendix), and can be digitally performed also in Alg. 5. Besides, Step 2 in Alg. 4 involves two matrix-matrix products, each with complexity $8X^{3}$ (see Appendix), leading to a complexity of $16X^{3}$. Thus, it is replaced by step 2 in Alg. 5, requiring the same complexity as Alg. 2, i.e., $6X^{2}$ real operations. In the “correction” phase, the computation of $\hat{\mathbf{x}}_{t\mid t}$ in Alg. 4 through Steps 3 and 4 requires two matrix-matrix products and a matrix inversion, leading to a complexity of $8(XY^{2}+X^{2}Y+Y^{3}/3)$, as for the LMMSE estimator analyzed in Remark 3. To reduce complexity, these steps are replaced by Steps 3, 4, and 5 in Alg. 5, having a respective complexity of $8X^{2}$ (matrix-vector product, see Appendix), $6XY$ (same as Alg. 1, see Remark 3), and $2X$ (sum of complex vectors, see Appendix). Finally, Step 5 in Alg. 4 requires two matrix-matrix products, and has complexity $8(X^{2}Y+X^{3})$. Thus, it is replaced by Steps 6 and 7 in Alg. 5 having lower complexity, i.e., $6XY$ (same as Alg. 2) and $4X^{2}$ (same as Alg. 3), respectively.
@@
A MiLAC can compute the Kalman filter with much lower computational complexity compared to conventional digital computing. Specifically, the computational complexity of Alg. 5 is given by adding the complexity of its Steps 1-7, i.e., $8X^{2}+6X^{2}+8X^{2}+6XY+2X+6XY+4X^{2}$, giving approximately $26X^{2}+12XY$ real operations. In contrast, the complexity of digitally computing the Kalman filter via Alg. 4 is given by adding the complexity of its Steps 1-5, namely $324X^{3}+16X^{2}Y+8XY^{2}+8Y^{3}/$ 3 real operations approximately. The complexity of computing the Kalman filter with a MiLAC and by digital computing are compared in Fig. 7, where we observe a gain of $1.1\times 10^{4}$ times when $X=Y=8192$. This decrease in complexity could have critical implications in the field of financial forecasting, where Kalman filters are extensively used to predict the trend of markets and assets with stringent latency requirements [43].
@@ VI Practical Implementation of a MiLAC with a Lossless Microwave Network
when $\bar{\mathbf{B}}$ is set as (60)-(63). Equivalently, we can introduce the permutation matrix $\boldsymbol{\Pi}\in\mathbb{C}^{2P\times 2P}$ as
@@
since $\boldsymbol{\Pi}^{T}\boldsymbol{\Pi}=\mathbf{I}_{2P}$. By recalling $\mathbf{v}=[\mathbf{v}_{1}^{T},\mathbf{v}_{2}^{T}]^{T}$ and $\tilde{\mathbf{u}}=[\mathbf{u}^{T},\mathbf{0}_{M\times 1}^{T}]^{T}$, (66) becomes
@@
where the term $\boldsymbol{\Pi}(j\bar{\mathbf{B}}/Y_{0}+\mathbf{I}_{2P})\boldsymbol{\Pi}^{T}$ can be rewritten as
@@
since $\boldsymbol{\Pi}\boldsymbol{\Pi}^{T}=\mathbf{I}_{2P}$ and $\bar{\mathbf{B}}$ is set as in (60)-(63). Substituting (68)-(70) into (67) and executing the products, (67) simplifies as
@@
which holds true given the relationship in (11)-(12), proving that a fully capable MiLAC as analyzed in Section III-A can be implemented with lossless microwave network. Note that a drawback of using lossless admittance components, which have purely imaginary values, is that the number of required components increases to $(2P)^{2}$, compared to the $P^{2}$ components needed to implement a MiLAC with complex admittance components.
@@ VII Conclusion
Analog computing is undergoing a revival, promising low-power and massively parallel computation capabilities for applications in signal processing and communications. In Part I of this paper, we introduce the concept of MiLAC as a generic analog computer that linearly operates with microwave signals, and analyze its application to signal processing. A MiLAC can generally be modeled as a multiport microwave network made of tunable impedance (or admittance) components. At the ports of such a microwave network, the input signals are applied and the output signals are read. By using rigorous microwave theory, we derive the expression of the output of a MiLAC as a function of its input and the values of its admittance components. We show that a MiLAC can efficiently compute the LMMSE estimator and invert an arbitrary matrix in the analog domain, with significantly reduced computational complexity over conventional digital computing. In particular, matrix inversion can be performed with complexity growing with the square of the matrix size, rather than the cube as with digital computing. The primitive operations computable with a MiLAC can serve as building blocks for implementing more complex algorithms, such as the Kalman filter, directly in the analog domain, achieving a substantial decrease in computational complexity. Finally, to facilitate the practical implementation of a MiLAC, we present a design based on lossless admittance components, avoiding losses and costly active RF components. In Part II of this paper, we discuss the application of MiLAC to wireless communications, showing that it can enable gigantic MIMO communications by performing computations and beamforming in the analog domain.
@@ VII-1 Scalar Product
Given two vectors $\mathbf{a}\in\mathbb{C}^{1\times N}$ and $\mathbf{b}\in\mathbb{C}^{N\times 1}$, the scalar product $\mathbf{a}\mathbf{b}$ requires $N-1$ complex additions and $N$ complex multiplications, since it is given by $[\mathbf{a}]_{1}[\mathbf{b}]_{1}+\ldots+[\mathbf{a}]_{N}[\mathbf{b}]_{N}$. Considering that complex additions and multiplications require two and six real operations, respectively, the complexity of the scalar product is approximately $8N$ real operations.
@@ VII-3 Matrix-Matrix Product
Given two matrices $\mathbf{A}\in\mathbb{C}^{M\times N}$ and $\mathbf{B}\in\mathbb{C}^{N\times L}$, the matrix-matrix product $\mathbf{A}\mathbf{B}$ can be expressed as $\mathbf{A}\mathbf{B}=[\mathbf{A}\mathbf{b}_{1},\ldots,\mathbf{A}\mathbf{b}_{L}]$, where $\mathbf{b}_{\ell}\in\mathbb{C}^{N\times 1}$ is the $\ell$th column of $\mathbf{B}$, for $\ell=1,\ldots,L$. Thus, it requires $L$ times the complexity of the matrix-vector product $\mathbf{A}\mathbf{b}$, which is approximately $8LMN$ real operations.
@@ VII-4 Matrix Inversion
Given an invertible matrix $\mathbf{A}\in\mathbb{C}^{N\times N}$, the algorithm commonly used to compute the matrix inverse $\mathbf{A}^{-1}$ is the Gauss’s method [44, Chapter 1]. Thus, we regard the complexity of matrix inversion as equal to the complexity of the Gauss’s method algorithm. It can be shown following [44, Chapter 1] that the Gauss’s method requires $N(N+1)/2$ divisions, $(2N^{3}+3N^{2}-5N)/6$ multiplications, and $(2N^{3}+3N^{2}-5N)/6$ subtractions. Consequently, the algorithm’s complexity is dominated by approximately $N^{3}/3$ multiplications and $N^{3}/3$ subtractions, which require a total of $38N^{3}/$ 3 real operations.

### v3

@@ V-B Computing the Kalman Filter with a MiLAC
The Kalman filter in Alg. 4 can be computed by performing the most computationally expensive steps in the analog domain through a MiLAC. This is shown in Alg. 5, which takes in input the matrix inverse $\mathbf{R}_{t-1\mid t-1}^{-1}$ and return the matrix inverse $\mathbf{R}_{t\mid t}^{-1}$, differently from Alg. 4, but without altering the logic of the Kalman filter. In the “prediction” phase, Step 1 in Alg. 4 has low computational complexity, i.e., $8X^{2}$ real operations (see Appendix), and can be digitally performed also in Alg. 5. Besides, Step 2 in Alg. 4 involves two matrix-matrix products, each with complexity $8X^{3}$ (see Appendix), leading to a complexity of $16X^{3}$. Thus, it is replaced by step 2 in Alg. 5, requiring the same complexity as Alg. 2, i.e., $6X^{2}$ real operations. In the “correction” phase, the computation of $\hat{\mathbf{x}}_{t\mid t}$ in Alg. 4 through Steps 3 and 4 requires two matrix-matrix products and a matrix inversion, leading to a complexity of $8(XY^{2}+X^{2}Y+Y^{3}/3)$, as for the LMMSE estimator analyzed in Remark 3. To reduce complexity, these steps are replaced by Steps 3, 4, and 5 in Alg. 5, having a respective complexity of $8X^{2}$ (matrix-vector product, see Appendix), $6XY$ (same as Alg. 1, see Remark 3), and $2X$ (sum of complex vectors, see Appendix). Finally, Step 5 in Alg. 4 requires two matrix-matrix products, and has complexity $8(X^{2}Y+X^{3})$. Thus, it is replaced by Steps 6 and 7 in Alg. 5 having lower complexity, i.e., $6XY$ (same as Alg. 2) and $4X^{2}$ (same as Alg. 3), respectively.
@@
A MiLAC can compute the Kalman filter with much lower computational complexity compared to conventional digital computing. Specifically, the computational complexity of Alg. 5 is given by adding the complexity of its Steps 1-7, i.e., $8X^{2}+6X^{2}+8X^{2}+6XY+2X+6XY+4X^{2}$, giving approximately $26X^{2}+12XY$ real operations. In contrast, the complexity of digitally computing the Kalman filter via Alg. 4 is given by adding the complexity of its Steps 1-5, namely $24X^{3}+16X^{2}Y+8XY^{2}+8Y^{3}/3$ real operations approximately. The complexity of computing the Kalman filter with a MiLAC and by digital computing are compared in Fig. 7, where we observe a gain of $1.1\times 10^{4}$ times when $X=Y=8192$. This decrease in complexity could have critical implications in the field of financial forecasting, where Kalman filters are extensively used to predict the trend of markets and assets with stringent latency requirements [54].
@@ VI Practical Implementation of a MiLAC with a Lossless Microwave Network
when $\bar{\mathbf{B}}$ is set as (60)-(63). Instead of proving (64), we can introduce the permutation matrix $\boldsymbol{\Pi}\in\mathbb{C}^{2P\times 2P}$ as
@@
which is equivalent to (64) since $\boldsymbol{\Pi}^{T}\boldsymbol{\Pi}=\mathbf{I}_{2P}$. By recalling $\mathbf{v}=[\mathbf{v}_{1}^{T},\mathbf{v}_{2}^{T}]^{T}$ and $\tilde{\mathbf{u}}=[\mathbf{u}^{T},\mathbf{0}_{M\times 1}^{T}]^{T}$, (66) becomes
@@
in which the term $\boldsymbol{\Pi}(j\bar{\mathbf{B}}/Y_{0}+\mathbf{I}_{2P})\boldsymbol{\Pi}^{T}$ can be rewritten as
@@
where in (68) we used $\boldsymbol{\Pi}\boldsymbol{\Pi}^{T}=\mathbf{I}_{2P}$ and in (69)-(70) we rewrote $\boldsymbol{\Pi}\bar{\mathbf{B}}\boldsymbol{\Pi}^{T}$ recalling that $\bar{\mathbf{B}}$ is set as in (60)-(63). Substituting (68)-(70) into (67) and executing the products, (67) simplifies as
@@
which holds true given the relationship in (11)-(12), proving that a fully capable MiLAC as analyzed in Section III-A can be implemented with lossless microwave network. Notably, a drawback of using lossless admittance components, which have purely imaginary values, is that the number of required components increases to $(2P)^{2}$, compared to the $P^{2}$ components needed to implement a MiLAC with complex admittance components.
@@ VII Conclusion
Analog computing is undergoing a revival, promising low-power and massively parallel computation capabilities for applications in signal processing and communications. In Part I of this paper, we introduce the concept of MiLAC as a generic analog computer that linearly operates with microwave signals, and analyze its application to signal processing. A MiLAC can generally be modeled as a multiport microwave network made of tunable impedance (or admittance) components. At the ports of such a microwave network, the input signals are applied and the output signals are read. By using rigorous microwave theory, we derive the expression of the output of a MiLAC as a function of its input and the values of its admittance components. We show that a MiLAC can efficiently compute the LMMSE estimator and invert an arbitrary matrix in the analog domain, with significantly reduced computational complexity over conventional digital computing. In particular, matrix inversion can be performed with complexity growing with the square of the matrix size, rather than the cube as with digital computing. The primitive operations computable with a MiLAC can serve as building blocks for implementing more complex algorithms, such as the Kalman filter, directly in the analog domain, achieving a substantial decrease in computational complexity. While we have shown how to efficiently compute certain common operations with MiLAC, future research could investigate whether other operations can also be efficiently computed in the analog domain with reconfigurable microwave networks. Finally, to facilitate the practical implementation of a MiLAC, we present a design based on lossless admittance components, avoiding losses and costly active RF components. Further future research could investigate the computational capabilities of microwave networks in the presence of hardware non-idealities, such as lossy interconnections. For instance, this can be done departing from previous literature on BD-RIS [55], where fully-connected microwave networks have been represented by separating the effects of the interconnections and the tunable components. Also assessing the power consumption of MiLAC and validating its practical benefits are interesting future research directions.
In Part II of this paper, we discuss the application of MiLAC to wireless communications, showing that it can enable gigantic MIMO communications by performing computations and beamforming in the analog domain.
@@ VII-1 Scalar Product
Given two vectors $\mathbf{a}\in\mathbb{C}^{1\times N}$ and $\mathbf{b}\in\mathbb{C}^{N\times 1}$, the scalar product $\mathbf{a}\mathbf{b}$ requires N1N-1 complex additions and $N$ complex multiplications, since it is given by $\mathbf{a}]_{1}[\mathbf{b}]_{1}+\ldots+[\mathbf{a}]_{N}[\mathbf{b}]_{N}$. Considering that complex additions and multiplications require two and six real operations, respectively, the complexity of the scalar product is approximately $8N$ real operations.
@@ VII-3 Matrix-Matrix Product
[…]
@@ VII-4 Matrix Inversion
Given an invertible matrix $\mathbf{A}\in\mathbb{C}^{N\times N}$, the algorithm commonly used to compute the matrix inverse $\mathbf{A}^{-1}$ is the Gauss’s method [56, Chapter 1]. Thus, we regard the complexity of matrix inversion as equal to the complexity of the Gauss’s method algorithm. It can be shown following [56, Chapter 1] that the Gauss’s method requires $N(N+1)/2$ divisions, $(2N^{3}+3N^{2}-5N)/6$ multiplications, and $(2N^{3}+3N^{2}-5N)/6$ subtractions. Consequently, the algorithm’s complexity is dominated by approximately $N^{3}/3$ multiplications and $N^{3}/3$ subtractions, which require a total of $8N^{3}/3$ real operations.